OpenStudy (anonymous):

What are the values of (-8)2 and -82

OpenStudy (calculusxy):

(-8)2 is just \[-8 \times 2\]

OpenStudy (calculusxy):

\[negative \times positive = negative\]

OpenStudy (anonymous):

these are the answers 64 and 64 16 and -16 64 and -64 16 and 16

OpenStudy (calculusxy):

Wait. So what does -82 supposed to mean?

OpenStudy (anonymous):

they want to know the values of them

OpenStudy (luffingsails):

Is it -8*2 ?

OpenStudy (anonymous):

no it is just -82 and (-8)squared

OpenStudy (calculusxy):

Ohh.. So i think that for -82 you multiply 8 times 2 for -8^2 = -8 x -8 = 64

OpenStudy (anonymous):

oh ok thank you!!!!

OpenStudy (calculusxy):

No problem :)

OpenStudy (calculusxy):

Wait.. -82 is -8 times 2

OpenStudy (calculusxy):

a typo

OpenStudy (anonymous):

ok tnx again

OpenStudy (calculusxy):

You're welcome

OpenStudy (anonymous):

so both of the values are 64 right

OpenStudy (calculusxy):

no. remember for -82 you are doing -8 x 2 = -16

OpenStudy (anonymous):

so -16 and 64

OpenStudy (calculusxy):

yeah that's what i think so

OpenStudy (anonymous):

ok